Convert 16.5 (base 16) to decimal. Show the result with four digits after the decimal point. Show work.
Answer #1

I hope this is what you want. If you still have any doubt please let me know. I have explained everything.

Convert 16.5 hexadecimal into decimal

To covert hexadecimal into decimal we will use following equation

If hexadecimal has digit (d3d2d1d0) than its equivalent decimal value is

Decimal= d3*163+ d2*162+ d1*161+ d0*160   

for digit after .(dot) it is divided by power of 16 (similar like binary to decimal)

Hence over here 16.5 hexadecimal is given

Decimal=1*161+6*160+5*16-1

                =1*16+6*1+5/16

                =16+6+0.3125

                =22.3125

In another way we first convert hexadecimal into binary and then convert binary into decimal

Hexadecimal 16.5 into binary

16.5= 0001 0110. 0101 (each hexadecimal digit represented using 4 bit binary)                  

To convert hexadecimal number into binary use the following table

Hexa decimal    

Binary

0

0000

1

0001

2

0010

3

0011

4

0100

5

0101

6

0110

7

0111

8

1000

9

1001

A

1010

B

1011

C

1100

D

1101

E

1110

F

1111

Now convert this 0001 0110. 0101 into decimal

Decimal= 1*24+0*23+1*22+1*21+0*20+ 0*2-1+1*2-2+0*2-3+1*2-4

                =1*16+0*8+1*4+1*2+0*1+(0/2)+(1/22)+(0/23)+(1/24)

                =16+0+4+2+0+0+(1/4)+0+(1/16)

=16+0+4+2+0+0+0.25+0+0.0625

                =22.3125